Solve simaltaneously for x and y: 2x-3y=-8 and 3x+4y=39

Answer by sachi(548)

2x-3y=-8 ----1
3x+4y=39 -----2
multiplying eqn 1 by 3 & eqn 2 by 2 & subtracting
6x-9y=-24
6x+8y=78
- - -
----------
-17y=-102
y=6
& x=[-8+3*6]/2=5
ans